OpenStudy (yellowlegoguy99):

Which of the following represents inverse variation? y equals three minus x y equals two divided by x y divided by x equals negative five y equals negative two-thirds x

OpenStudy (yellowlegoguy99):

@Jack1

OpenStudy (yellowlegoguy99):

Medal to who answers best! C:

OpenStudy (jack1):

inverse variation is inversly proportional so eg doubling x causes y to halve. which one fits that scenario?

OpenStudy (jack1):

so opposite effect, if you do ___ to one opposite happens to the other

OpenStudy (jack1):

a) y equals three minus x so equation: y = 3 - x so if y = 10, and we double y, what happens to x y = 3 - x 10 = 3 - x x = -7 so double y 2y = 2(3 - x) 20 = 6 - 2x -14 = 2x so y is directly proportional to x in this doubling y doubled x

OpenStudy (debbieg):

Another way to think of it: If you can write y=k/x, then that is inverse variation. If you can write y=kx then that is direct variation

OpenStudy (anonymous):

inverse variation is when on increasing one variable the value of the other decreases. like y = k/x